# Differentiable and accelerated spherical harmonic and Wigner transforms

### v1

@@ 5 Gradients of spherical transforms
One of the primary goals of this work is not only to provide computational approaches for spherical transforms that can be deployed on hardware accelerators but that also facilitate the efficient computation of gradients. Gradients are essential for differentiable programming applications. For example, machine learning models on the sphere Cohen et al. (2018); Kondor et al. (2018); Cobb et al. (2021); McEwen et al. (2022) often require spherical transforms that are differentiable so that the models may be trained by gradient-based optimisation algorithms. Differentiable physical models are also required for hybrid data-driven and model-based approaches; in many cases an underlying component of such physical models includes spherical transforms, for example in cosmology (Dodelson, 2003) and seismology (Dahlen and Tromp, 1998).
@@ 5.1 Automatic differentiation (AD) modes
Modern automatic differentiation (AD) predominately relies on one of two gradient propagation schemes: forward- and reverse-mode. Forward-mode AD applies the chain rule sequentially to each primitive operation in a forward (primal) trace of a function, accumulating the primal and derivative (tangent) variables during a single forward pass Wengert (1964). Conversely, in reverse-mode AD forward and reverse passes are performed. During the forward pass primal variables are computed and propagated, with intermediate values stored in memory for use in the following reverse pass. In the reverse pass the derivatives (cotangents) with respect to intermediate values are propagated backwards Baydin et al. (2018). Reverse mode AD applied to training neural networks with a scalar cost function is commonly referred to as back-propagation Rumelhart et al. (1986). Forward- and reverse-mode AD provide efficient ways to compute, respectively, Jacobian-vector products and vector-Jacobian (transposed Jacobian-vector) products for the traced function.
The choice of which mode is most efficient is determined by the problem at hand. Consider the case in which we wish to compute the derivative of a function $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$. A full column of the Jacobian can be computed in a single pass of forward-mode AD, whereas a full row of the Jacobian can be computed in a single pass of reverse-mode AD Baydin et al. (2018). Consequently, when $n\ll m$ forward-mode AD is most efficient, whereas when $m\ll n$ reverse-mode is most efficient. However, note that intermediate values must be stored for reverse-mode AD, increasing memory requirements, which is not the case for forward-mode AD.
@@ 5.2 Automatic versus manual differentiation
By leveraging an AD framework, such as JAX, one can compute JVPs and VJPs simply through the framework. This approach avoids the need for any gradient-related implementation, thus eliminating complexity and possibilities for errors. However, as discussed above, reverse-mode AD induces a memory overhead that in the worst case can grow in proportion to the number of operations of the traced function Baydin et al. (2018). Furthermore, for machine learning applications, reverse-mode AD is the mode of interest since it is most efficient for gradient-based optimisation of scalar loss functions that depend on a very large number of input parameters. For the spherical transforms considered in this article, we indeed observed that the computation of VJPs purely by reverse-mode AD exhibited a very high memory overhead (as discussed further below).
An alternative to automatic differentiation is manual differentiation, where explicit expressions for JVPs and VJPs are determined analytically and then implemented directly. For linear transforms, such as the spherical transforms considered herein, JVPs and VPJs take relatively straightforward forms, where the JVP is simply the transform itself and the VJP is closely related to the adjoint transform. Algorithms to compute adjoints of spherical transforms efficiently have been constructed previously McEwen et al. (2013); Wallis et al. (2017); Price et al. (2021). Such a direct implementation of manual derivatives requires bespoke implementation, which can be cumbersome (especially to support the subtleties of different spherical samplings) and thus can be prone to errors. Nevertheless, a forward pass to compute VJPs is not required, eliminating the memory overhead that is otherwise necessary.
@@ 5.3.1 Forward spin spherical harmonic transform
Adjoint operators of transforms are commonly required (e.g. when solving inverse problems posed as variational regularisation problems by convex optimisation techniques; McEwen et al. (2013); Wallis et al. (2017); Price et al. (2021)). The VJP is closely related to the adjoint operator of the transform up to complex conjugation. For completeness we note the adjoint of the forward spherical harmonic transforms $\mathcal{F}$ can be represented as
@@ 5.4 Hybrid automatic and manual differentiable spherical transforms
We therefore consider a hybrid approach. We adopt manual differentiation for the core internal components of spherical transforms related to recursions but adopt AD for surrounding components of the transforms that deal with the subtleties of spherical samplings. In this manner we avoid the prohibitive memory overhead of a pure AD approach, while also avoiding the subtleties of cumbersome direct implementations of a manual differentiation approach.
@@ 6 Software and evaluation
A variety of differentiable programming and machine learning ecosystems exist in Python that could be considered for the implementation of our software, e.g. TensorFlow and PyTorch, both of which exhibit a mature collection of Python APIs. However, over time these libraries have become somewhat rigid due to the nature of heavily templated object oriented software focused on machine learning and hence the development of other bespoke functionality is hindered by, for example, significant boilerplate code. Consequently, we choose to develop our software in JAX, which is a functional differentiable programming Python framework recently developed by Google Bradbury et al. (2018). JAX is particularly well-suited to scientific software development, where flexibility and rapid development are highly desirable.
@@ 6.1 Sampling agnostic transforms
As discussed, our software is largely agnostic to the choice of spherical sampling considered. We simply require an isolatitudinal sampling, which most popular spherical samplings follow due to the considerable computational savings that it affords through a separation of variables, and corresponding quadrature weights. As a starting point Driscoll & Healy Driscoll and Healy (1994), McEwen & Wiaux (both MW and MWSS) McEwen and Wiaux (2011); Daducci et al. (2011); Ocampo et al. (2023), and HEALPix Górski et al. (2005) samplings are supported. To simplify further open-sourced contributions, our code is designed with integration of future sampling schemes in mind, should the need arise. It is worth noting that once sample locations and associated quadrature weights are defined, all downstream functionality should work seamlessly.
@@ 6.3 Benchmarking
We present comprehensive benchmarking of the core transforms provided by our S2FFT software. First, we quantify the precision of our transforms by computing the round-trip error and find errors to be at the level of numerical precision for the sampling theorems considered. Second, we compute the average wall-clock computational speed of transforms, in both single and multiple GPU scenarios, for both our on-the-fly algorithms (computing Wigner $d$-functions on-the-fly) and precompute algorithm (computing and storing Wigner $d$-functions a priori). We compare computational time to the SSHT [^4] https://github.com/astro-informatics/ssht McEwen and Wiaux (2011) and SO3555https://github.com/astro-informatics/so3 McEwen et al. (2015) codes that compute spherical harmonic and Wigner transforms, respectively, for which underlying algorithms are implemented in C. An extensive comparison against a wide variety of existing spherical transform software codes is beyond the scope of this study (furthermore, while there are numerous spherical harmonic transforms codes there are relatively few codes capable of also computing Wigner transforms). In any case, comparison against other codes can be inferred from contemporary papers. In addition, all algorithms implemented in S2FFT are validated through extensive unit and regression tests.
@@ 6.3.1 Precision
Natively, our software supports both 32-bit and 64-bit floating point precision, however due to the unstable nature of three term recursions Gautschi (1967) as outlined in Sec. 3 it is strongly recommended to adopt 64-bit precision. To mimic the most likely use-case, in our benchmarking we default to 64-bit precision, which is provided by JAX as an extended data type. Both Driscoll & Healy sampling Driscoll and Healy (1994) and both McEwen & Wiaux sampling schemes McEwen and Wiaux (2011) afford sampling theorems on the sphere and therefore transforms based on such samplings are theoretically exact to machine precision.
@@
In the era of high precision science, access to theoretically exact transforms is of great importance BICEP2 Collaboration (2014); Planck Collaboration I (2020); Pires et al. (2020); Amaro-Seoane et al. (2017). The best way to perform high precision science is to avoid introducing numerical errors in the first place. For other applications one may wish to adopt 32-bit floating point precision, which can further accelerate compute and will half memory overhead. Such applications may include geometric machine learning on the sphere Bronstein et al. (2021); Cohen et al. (2018); Kondor et al. (2018); Cobb et al. (2021), where 32-bit precision (or lower) is often adopted. When adopting 32-bit precision one may expect to recover commensurate errors, i.e. errors beginning at approximately $\mathcal{O}(10^{-8})$ rather than $\mathcal{O}(10^{-16})$.
@@
L
MW
MWSS
DH
MW
MWSS
DH
$8$
$3.6\times 10^{-16}$
$1.7\times 10^{-16}$
$5.1\times 10^{-16}$
$4.6\times 10^{-16}$
$4.3\times 10^{-16}$
$4.3\times 10^{-16}$
$16$
$3.7\times 10^{-16}$
$2.7\times 10^{-16}$
$6.3\times 10^{-16}$
$5.4\times 10^{-16}$
$4.5\times 10^{-16}$
$4.5\times 10^{-16}$
$32$
$7.5\times 10^{-16}$
$6.3\times 10^{-16}$
$3.5\times 10^{-16}$
$7.3\times 10^{-16}$
$7.2\times 10^{-16}$
$7.2\times 10^{-16}$
$64$
$1.2\times 10^{-15}$
$1.1\times 10^{-15}$
$6.7\times 10^{-16}$
$1.2\times 10^{-15}$
$1.2\times 10^{-15}$
$1.2\times 10^{-15}$
$128$
$2.3\times 10^{-15}$
$2.3\times 10^{-15}$
$1.3\times 10^{-15}$
$2.5\times 10^{-15}$
$2.4\times 10^{-15}$
$2.4\times 10^{-15}$
$256$
$4.7\times 10^{-15}$
$5.0\times 10^{-15}$
$2.6\times 10^{-15}$
$4.7\times 10^{-15}$
$4.7\times 10^{-15}$
$4.7\times 10^{-15}$
$512$
$1.0\times 10^{-14}$
$9.8\times 10^{-15}$
$4.6\times 10^{-15}$
$9.8\times 10^{-15}$
$9.7\times 10^{-15}$
$8.9\times 10^{-15}$
$1024$
$1.9\times 10^{-14}$
$1.9\times 10^{-14}$
$9.3\times 10^{-15}$
$1.7\times 10^{-14}$
$1.5\times 10^{-14}$
$1.2\times 10^{-14}$
$2048$
$3.7\times 10^{-14}$
$3.8\times 10^{-14}$
$1.9\times 10^{-14}$
$4096$
$7.5\times 10^{-14}$
$7.7\times 10^{-14}$
$3.8\times 10^{-14}$
$8192$
$1.5\times 10^{-13}$
$1.5\times 10^{-13}$
$8.3\times 10^{-14}$
@@
L
MW
MWSS
MW
MWSS
$8$
$1.6\times 10^{-15}$
$1.3\times 10^{-15}$
$1.3\times 10^{-15}$
$1.2\times 10^{-15}$
$16$
$1.2\times 10^{-15}$
$1.0\times 10^{-15}$
$1.1\times 10^{-15}$
$1.1\times 10^{-15}$
$32$
$1.3\times 10^{-15}$
$1.2\times 10^{-15}$
$1.5\times 10^{-15}$
$1.2\times 10^{-15}$
$64$
$1.5\times 10^{-15}$
$1.4\times 10^{-15}$
$1.8\times 10^{-15}$
$1.4\times 10^{-15}$
$128$
$2.2\times 10^{-15}$
$2.0\times 10^{-15}$
$3.6\times 10^{-15}$
$2.0\times 10^{-15}$
$256$
$2.9\times 10^{-15}$
$2.8\times 10^{-15}$
$4.2\times 10^{-15}$
$2.9\times 10^{-15}$
$512$
$4.2\times 10^{-15}$
$4.3\times 10^{-15}$
$6.4\times 10^{-15}$
$4.2\times 10^{-15}$
$1024$
$5.7\times 10^{-15}$
$5.9\times 10^{-15}$
$2048$
$8.1\times 10^{-15}$
$8.2\times 10^{-15}$
$4096$
$1.2\times 10^{-14}$
$1.3\times 10^{-14}$
@@
Fig. 6: Numerical benchmarking results for the spherical harmonic transforms implemented in S2FFT compared to the C implementation of SSHT. We consider samplings of the sphere that afford a sampling theorem and hence exact spherical harmonic transforms. S2FFT provides multiple operational modes: a precompute approach with faster compute at $\mathcal{O}(L^{3})$ memory overhead and an on-the-fly transform with negligible memory overhead (see Sec. 4). Top: Average compute time required for a spherical harmonic transform for each operational modality. Middle: Relative acceleration of a given modality with respect to their SSHT counterpart. Bottom: Averaged round-trip numerical error. In each case we recover machine precision. We also empirically observe $\mathcal{O}(L)$ error scaling (in line with previous work McEwen and Wiaux (2011)). The silver and gold stars represent the performance one may expect to realise should one distribute our on-the-fly algorithms across 100 and 1000 GPUs respectively.
@@ 6.3.2 Computational speed
$L$
Time SSHT
Time S2FFT
Speed-Up 1 $\times$ GPU
Speed-Up 3 $\times$ GPU
3 $\times$ GPU Ratio
Time S2FFT
Speed-Up 1 $\times$ GPU
$8$
$8.8\times 10^{-2}$
$3.3\times 10^{-1}$
0.26
0.07
0.27
$4.4\times 10^{-2}$
2.03
$16$
$4.5\times 10^{-1}$
$9.2\times 10^{-1}$
0.49
0.23
0.47
$4.1\times 10^{-2}$
11.0
$32$
$2.2\times 10^{0}$
$2.0\times 10^{0}$
1.08
0.73
0.68
$4.0\times 10^{-2}$
53.9
$64$
$1.4\times 10^{1}$
$3.6\times 10^{0}$
3.87
2.83
0.73
$7.6\times 10^{-2}$
184
$128$
$8.6\times 10^{1}$
$7.8\times 10^{0}$
11.0
8.74
0.80
$2.9\times 10^{-1}$
299
$256$
$6.2\times 10^{2}$
$2.4\times 10^{1}$
26.1
25.8
0.99
$1.3\times 10^{0}$
475
$512$
$4.6\times 10^{3}$
$1.2\times 10^{2}$
38.1
64.1
1.68
$7.1\times 10^{0}$
646
$1024$
$3.6\times 10^{4}$
$8.6\times 10^{2}$
41.5
103
2.48
$6.0\times 10^{1}$
592
$2048$
$2.9\times 10^{5}$
$7.5\times 10^{3}$
39.0
104
2.67
$4096$
$2.4\times 10^{6}$
$5.1\times 10^{4}$
46.4
135
2.91
$8192$
$1.9\times 10^{7}$
$4.1\times 10^{5}$
47.0
139
2.96
$8$
$5.0\times 10^{-2}$
$3.3\times 10^{-1}$
0.15
0.04
0.27
$4.1\times 10^{-2}$
1.20
$16$
$1.5\times 10^{-1}$
$9.0\times 10^{-1}$
0.16
0.07
0.44
$4.0\times 10^{-2}$
3.63
$32$
$3.8\times 10^{-1}$
$1.9\times 10^{0}$
0.20
0.13
0.65
$4.3\times 10^{-2}$
8.94
$64$
$3.2\times 10^{0}$
$3.6\times 10^{0}$
0.88
0.62
0.70
$5.2\times 10^{-2}$
60.5
$128$
$1.3\times 10^{1}$
$7.3\times 10^{0}$
1.80
1.42
0.79
$1.6\times 10^{-1}$
80.6
$256$
$1.1\times 10^{2}$
$1.7\times 10^{1}$
6.32
4.95
0.78
$6.7\times 10^{-1}$
163
$512$
$6.6\times 10^{2}$
$7.5\times 10^{1}$
8.85
11.4
1.29
$3.6\times 10^{0}$
184
$1024$
$6.4\times 10^{3}$
$4.3\times 10^{2}$
14.8
32.9
2.22
$3.3\times 10^{1}$
196
$2048$
$7.2\times 10^{4}$
$3.9\times 10^{3}$
18.6
49.7
2.67
$4096$
$1.1\times 10^{6}$
$2.4\times 10^{4}$
46.7
134
2.87
$8192$
$9.1\times 10^{6}$
$2.4\times 10^{5}$
37.4
111
2.97
$8$
$4.8\times 10^{-2}$
$4.3\times 10^{-1}$
0.11
0.03
0.27
$4.6\times 10^{-2}$
1.05
$16$
$1.2\times 10^{-1}$
$8.6\times 10^{-1}$
0.14
0.06
0.43
$4.5\times 10^{-2}$
2.60
$32$
$5.5\times 10^{-1}$
$1.8\times 10^{0}$
0.30
0.19
0.63
$4.7\times 10^{-2}$
11.8
$64$
$3.8\times 10^{0}$
$3.5\times 10^{0}$
1.09
0.79
0.72
$5.0\times 10^{-2}$
76.9
$128$
$2.7\times 10^{1}$
$7.2\times 10^{0}$
3.82
3.06
0.80
$1.5\times 10^{-1}$
180
$256$
$2.1\times 10^{2}$
$1.7\times 10^{1}$
12.4
9.80
0.79
$6.6\times 10^{-1}$
327
$512$
$1.6\times 10^{3}$
$6.8\times 10^{1}$
24.0
29.7
1.24
$3.6\times 10^{0}$
459
$1024$
$1.3\times 10^{4}$
$4.5\times 10^{2}$
28.7
64.2
2.24
$3.3\times 10^{1}$
400
$2048$
$1.5\times 10^{5}$
$3.8\times 10^{3}$
39.6
102
2.58
$4096$
$2.2\times 10^{6}$
$2.6\times 10^{4}$
83.6
234
2.80
$8192$
$1.8\times 10^{7}$
$2.6\times 10^{5}$
68.7
197
2.87

### v2

@@ 5 Gradients of spherical transforms
One of the primary goals of this work is not only to provide computational approaches for spherical transforms that can be deployed on hardware accelerators but that also facilitate the efficient computation of gradients. Gradients are essential for differentiable programming applications. For example, machine learning models on the sphere [66, 4, 32, 33] often require spherical transforms that are differentiable so that the models may be trained by gradient-based optimisation algorithms. Differentiable physical models are also required for hybrid data-driven and model-based approaches; in many cases an underlying component of such physical models includes spherical transforms, for example in cosmology [25] and seismology [26].
@@ 5.1 Automatic differentiation (AD) modes
Modern automatic differentiation (AD) predominately relies on one of two gradient propagation schemes: forward- and reverse-mode. Forward-mode AD applies the chain rule sequentially to each primitive operation in a forward (primal) trace of a function, accumulating the primal and derivative (tangent) variables during a single forward pass [67]. Conversely, in reverse-mode AD forward and reverse passes are performed. During the forward pass primal variables are computed and propagated, with intermediate values stored in memory for use in the following reverse pass. In the reverse pass the derivatives (cotangents) with respect to intermediate values are propagated backwards [68]. Reverse mode AD applied to training neural networks with a scalar cost function is commonly referred to as back-propagation [69]. Forward- and reverse-mode AD provide efficient ways to compute, respectively, Jacobian-vector products and vector-Jacobian (transposed Jacobian-vector) products for the traced function.
The choice of which mode is most efficient is determined by the problem at hand. Consider the case in which we wish to compute the derivative of a function $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$. A full column of the Jacobian can be computed in a single pass of forward-mode AD, whereas a full row of the Jacobian can be computed in a single pass of reverse-mode AD [68]. Consequently, when $n\ll m$ forward-mode AD is most efficient, whereas when $m\ll n$ reverse-mode is most efficient. However, note that intermediate values must be stored for reverse-mode AD, increasing memory requirements, which is not the case for forward-mode AD.
@@ 5.2 Automatic versus manual differentiation
By leveraging an AD framework, such as JAX, one can compute JVPs and VJPs simply through the framework. This approach avoids the need for any gradient-related implementation, thus eliminating complexity and possibilities for errors. However, as discussed above, reverse-mode AD induces a memory overhead that in the worst case can grow in proportion to the number of operations of the traced function [68]. Furthermore, for machine learning applications, reverse-mode AD is the mode of interest since it is most efficient for gradient-based optimisation of scalar loss functions that depend on a very large number of input parameters. For the spherical transforms considered in this article, we indeed observed that the computation of VJPs purely by reverse-mode AD exhibited a very high memory overhead (as discussed further below).
An alternative to automatic differentiation is manual differentiation, where explicit expressions for JVPs and VJPs are determined analytically and then implemented directly. For linear transforms, such as the spherical transforms considered herein, JVPs and VPJs take relatively straightforward forms, where the JVP is simply the transform itself and the VJP is closely related to the adjoint transform. Algorithms to compute adjoints of spherical transforms efficiently have been constructed previously [70, 71, 72]. Such a direct implementation of manual derivatives requires bespoke implementation, which can be cumbersome (especially to support the subtleties of different spherical samplings) and thus can be prone to errors. Nevertheless, a forward pass to compute VJPs is not required, eliminating the memory overhead that is otherwise necessary.
@@ 5.3.1 Forward spin spherical harmonic transform
Adjoint operators of transforms are commonly required (e.g. when solving inverse problems posed as variational regularisation problems by convex optimisation techniques; [70, 71, 72]). The VJP is closely related to the adjoint operator of the transform up to complex conjugation. For completeness we note the adjoint of the forward spherical harmonic transforms $\mathcal{F}$ can be represented as
@@ 5.4 Hybrid automatic and manual differentiable spherical transforms
We therefore consider a hybrid approach. We adopt manual differentiation for the core internal components of spherical transforms related to recursions but adopt AD for surrounding components of the transforms that deal with the subtleties of spherical samplings. In this manner we avoid the prohibitive memory overhead of a pure AD approach, while also avoiding the subtleties of cumbersome direct implementations of a manual differentiation approach. A similar hybrid gradient approach was recently considered in Li et al. [73] to mitigate a similar JAX memory issue whilst differentiating through cosmological simulations.
@@ 6 Software and evaluation
A variety of differentiable programming and machine learning ecosystems exist in Python that could be considered for the implementation of our software, e.g. TensorFlow and PyTorch, both of which exhibit a mature collection of Python APIs. However, over time these libraries have become somewhat rigid due to the nature of heavily templated object oriented software focused on machine learning and hence the development of other bespoke functionality is hindered by, for example, significant boilerplate code. Consequently, we choose to develop our software in JAX, which is a functional differentiable programming Python framework recently developed by Google [56]. JAX is particularly well-suited to scientific software development, where flexibility and rapid development are highly desirable.
@@ 6.1 Sampling agnostic transforms
As discussed, our software is largely agnostic to the choice of spherical sampling considered. We simply require an isolatitudinal sampling, which most popular spherical samplings follow due to the considerable computational savings that it affords through a separation of variables, and corresponding quadrature weights. As a starting point Driscoll & Healy [40], McEwen & Wiaux (both MW and MWSS) [41, 6, 64], Gauss-Lengendre [e.g. 54, 41], and HEALPix [43] samplings are supported. To simplify further open-sourced contributions, our code is designed with integration of future sampling schemes in mind, should the need arise. It is worth noting that once sample locations and associated quadrature weights are defined, all downstream functionality should work seamlessly.
@@ 6.3 Benchmarking
We present comprehensive benchmarking of the core transforms provided by our S2FFT software. First, we quantify the precision of our transforms by computing the round-trip error and find errors to be at the level of numerical precision for the sampling theorems considered. Second, we compute the average wall-clock computational speed of transforms, in both single and multiple GPU scenarios, for both our on-the-fly algorithms (computing Wigner $d$-functions on-the-fly) and precompute algorithm (computing and storing Wigner $d$-functions a priori). We compare computational time to the SSHT [^4] https://github.com/astro-informatics/ssht [41] and SO3555https://github.com/astro-informatics/so3 [42] codes that compute spherical harmonic and Wigner transforms, respectively, for which underlying algorithms are implemented in C. An extensive comparison against a wide variety of existing spherical transform software codes is beyond the scope of this study (furthermore, while there are numerous spherical harmonic transforms codes there are relatively few codes capable of also computing Wigner transforms). In any case, comparison against other codes can be inferred from contemporary papers [see e.g. 41, 61]. In addition, all algorithms implemented in S2FFT are validated through extensive unit and regression tests.
@@ 6.3.1 Precision
Natively, our software supports both 32-bit and 64-bit floating point precision, however due to the unstable nature of three term recursions [59] as outlined in Sec. 3 it is strongly recommended to adopt 64-bit precision. To mimic the most likely use-case, in our benchmarking we default to 64-bit precision, which is provided by JAX as an extended data type. Both Driscoll & Healy sampling [40] and both McEwen & Wiaux sampling schemes [41] afford sampling theorems on the sphere and therefore transforms based on such samplings are theoretically exact to machine precision.
@@
In the era of high precision science, access to theoretically exact transforms is of great importance [74, 20, 21, 24]. The best way to perform high precision science is to avoid introducing numerical errors in the first place. For other applications one may wish to adopt 32-bit floating point precision, which can further accelerate compute and will half memory overhead. Such applications may include geometric machine learning on the sphere [75, 30, 4, 32], where 32-bit precision (or lower) is often adopted.
@@
L MW MWSS DH MW MWSS DH
$8$ $3.6\times 10^{-16}$ $1.7\times 10^{-16}$ $5.1\times 10^{-16}$ $4.6\times 10^{-16}$ $4.3\times 10^{-16}$ $4.3\times 10^{-16}$
$16$ $3.7\times 10^{-16}$ $2.7\times 10^{-16}$ $6.3\times 10^{-16}$ $5.4\times 10^{-16}$ $4.5\times 10^{-16}$ $4.5\times 10^{-16}$
$32$ $7.5\times 10^{-16}$ $6.3\times 10^{-16}$ $3.5\times 10^{-16}$ $7.3\times 10^{-16}$ $7.2\times 10^{-16}$ $7.2\times 10^{-16}$
$64$ $1.2\times 10^{-15}$ $1.1\times 10^{-15}$ $6.7\times 10^{-16}$ $1.2\times 10^{-15}$ $1.2\times 10^{-15}$ $1.2\times 10^{-15}$
$128$ $2.3\times 10^{-15}$ $2.3\times 10^{-15}$ $1.3\times 10^{-15}$ $2.5\times 10^{-15}$ $2.4\times 10^{-15}$ $2.4\times 10^{-15}$
$256$ $4.7\times 10^{-15}$ $5.0\times 10^{-15}$ $2.6\times 10^{-15}$ $4.7\times 10^{-15}$ $4.7\times 10^{-15}$ $4.7\times 10^{-15}$
$512$ $1.0\times 10^{-14}$ $9.8\times 10^{-15}$ $4.6\times 10^{-15}$ $9.8\times 10^{-15}$ $9.7\times 10^{-15}$ $8.9\times 10^{-15}$
$1024$ $1.9\times 10^{-14}$ $1.9\times 10^{-14}$ $9.3\times 10^{-15}$ $1.7\times 10^{-14}$ $1.5\times 10^{-14}$ $1.2\times 10^{-14}$
$2048$ $3.7\times 10^{-14}$ $3.8\times 10^{-14}$ $1.9\times 10^{-14}$
$4096$ $7.5\times 10^{-14}$ $7.7\times 10^{-14}$ $3.8\times 10^{-14}$
$8192$ $1.5\times 10^{-13}$ $1.5\times 10^{-13}$ $8.3\times 10^{-14}$
@@
L MW MWSS MW MWSS
$8$ $1.6\times 10^{-15}$ $1.3\times 10^{-15}$ $1.3\times 10^{-15}$ $1.2\times 10^{-15}$
$16$ $1.2\times 10^{-15}$ $1.0\times 10^{-15}$ $1.1\times 10^{-15}$ $1.1\times 10^{-15}$
$32$ $1.3\times 10^{-15}$ $1.2\times 10^{-15}$ $1.5\times 10^{-15}$ $1.2\times 10^{-15}$
$64$ $1.5\times 10^{-15}$ $1.4\times 10^{-15}$ $1.8\times 10^{-15}$ $1.4\times 10^{-15}$
$128$ $2.2\times 10^{-15}$ $2.0\times 10^{-15}$ $3.6\times 10^{-15}$ $2.0\times 10^{-15}$
$256$ $2.9\times 10^{-15}$ $2.8\times 10^{-15}$ $4.2\times 10^{-15}$ $2.9\times 10^{-15}$
$512$ $4.2\times 10^{-15}$ $4.3\times 10^{-15}$ $6.4\times 10^{-15}$ $4.2\times 10^{-15}$
$1024$ $5.7\times 10^{-15}$ $5.9\times 10^{-15}$
$2048$ $8.1\times 10^{-15}$ $8.2\times 10^{-15}$
$4096$ $1.2\times 10^{-14}$ $1.3\times 10^{-14}$
@@
Fig. 6: Numerical benchmarking results for the spherical harmonic transforms implemented in S2FFT compared to the C implementation of SSHT. We consider samplings of the sphere that afford a sampling theorem and hence exact spherical harmonic transforms. S2FFT provides multiple operational modes: a precompute approach with faster compute at $\mathcal{O}(L^{3})$ memory overhead and an on-the-fly transform with negligible memory overhead (see Sec. 4). Top: Average compute time required for a spherical harmonic transform for each operational modality. Middle: Relative acceleration of a given modality with respect to their SSHT counterpart. Bottom: Averaged round-trip numerical error. In each case we recover machine precision. We also empirically observe $\mathcal{O}(L)$ error scaling (in line with previous work [41]). The silver and gold stars represent the performance one may expect to realise should one distribute our on-the-fly algorithms across 100 and 1000 GPUs respectively.
@@ 6.3.2 Computational speed
$L$ Time SSHT Time S2FFT Speed-Up 1 $\times$ GPU Speed-Up 3 $\times$ GPU 3 $\times$ GPU Ratio Time S2FFT Speed-Up 1 $\times$ GPU
$8$ $8.8\times 10^{-2}$ $3.3\times 10^{-1}$ 0.26 0.07 0.27 $4.4\times 10^{-2}$ 2.03
$16$ $4.5\times 10^{-1}$ $9.2\times 10^{-1}$ 0.49 0.23 0.47 $4.1\times 10^{-2}$ 11.0
$32$ $2.2\times 10^{0}$ $2.0\times 10^{0}$ 1.08 0.73 0.68 $4.0\times 10^{-2}$ 53.9
$64$ $1.4\times 10^{1}$ $3.6\times 10^{0}$ 3.87 2.83 0.73 $7.6\times 10^{-2}$ 184
$128$ $8.6\times 10^{1}$ $7.8\times 10^{0}$ 11.0 8.74 0.80 $2.9\times 10^{-1}$ 299
$256$ $6.2\times 10^{2}$ $2.4\times 10^{1}$ 26.1 25.8 0.99 $1.3\times 10^{0}$ 475
$512$ $4.6\times 10^{3}$ $1.2\times 10^{2}$ 38.1 64.1 1.68 $7.1\times 10^{0}$ 646
$1024$ $3.6\times 10^{4}$ $8.6\times 10^{2}$ 41.5 103 2.48 $6.0\times 10^{1}$ 592
$2048$ $2.9\times 10^{5}$ $7.5\times 10^{3}$ 39.0 104 2.67
$4096$ $2.4\times 10^{6}$ $5.1\times 10^{4}$ 46.4 135 2.91
$8192$ $1.9\times 10^{7}$ $4.1\times 10^{5}$ 47.0 139 2.96
$8$ $5.0\times 10^{-2}$ $3.3\times 10^{-1}$ 0.15 0.04 0.27 $4.1\times 10^{-2}$ 1.20
$16$ $1.5\times 10^{-1}$ $9.0\times 10^{-1}$ 0.16 0.07 0.44 $4.0\times 10^{-2}$ 3.63
$32$ $3.8\times 10^{-1}$ $1.9\times 10^{0}$ 0.20 0.13 0.65 $4.3\times 10^{-2}$ 8.94
$64$ $3.2\times 10^{0}$ $3.6\times 10^{0}$ 0.88 0.62 0.70 $5.2\times 10^{-2}$ 60.5
$128$ $1.3\times 10^{1}$ $7.3\times 10^{0}$ 1.80 1.42 0.79 $1.6\times 10^{-1}$ 80.6
$256$ $1.1\times 10^{2}$ $1.7\times 10^{1}$ 6.32 4.95 0.78 $6.7\times 10^{-1}$ 163
$512$ $6.6\times 10^{2}$ $7.5\times 10^{1}$ 8.85 11.4 1.29 $3.6\times 10^{0}$ 184
$1024$ $6.4\times 10^{3}$ $4.3\times 10^{2}$ 14.8 32.9 2.22 $3.3\times 10^{1}$ 196
$2048$ $7.2\times 10^{4}$ $3.9\times 10^{3}$ 18.6 49.7 2.67
$4096$ $1.1\times 10^{6}$ $2.4\times 10^{4}$ 46.7 134 2.87
$8192$ $9.1\times 10^{6}$ $2.4\times 10^{5}$ 37.4 111 2.97
$8$ $4.8\times 10^{-2}$ $4.3\times 10^{-1}$ 0.11 0.03 0.27 $4.6\times 10^{-2}$ 1.05
$16$ $1.2\times 10^{-1}$ $8.6\times 10^{-1}$ 0.14 0.06 0.43 $4.5\times 10^{-2}$ 2.60
$32$ $5.5\times 10^{-1}$ $1.8\times 10^{0}$ 0.30 0.19 0.63 $4.7\times 10^{-2}$ 11.8
$64$ $3.8\times 10^{0}$ $3.5\times 10^{0}$ 1.09 0.79 0.72 $5.0\times 10^{-2}$ 76.9
$128$ $2.7\times 10^{1}$ $7.2\times 10^{0}$ 3.82 3.06 0.80 $1.5\times 10^{-1}$ 180
$256$ $2.1\times 10^{2}$ $1.7\times 10^{1}$ 12.4 9.80 0.79 $6.6\times 10^{-1}$ 327
$512$ $1.6\times 10^{3}$ $6.8\times 10^{1}$ 24.0 29.7 1.24 $3.6\times 10^{0}$ 459
$1024$ $1.3\times 10^{4}$ $4.5\times 10^{2}$ 28.7 64.2 2.24 $3.3\times 10^{1}$ 400
$2048$ $1.5\times 10^{5}$ $3.8\times 10^{3}$ 39.6 102 2.58
$4096$ $2.2\times 10^{6}$ $2.6\times 10^{4}$ 83.6 234 2.80
$8192$ $1.8\times 10^{7}$ $2.6\times 10^{5}$ 68.7 197 2.87